OpenStudy (diamondboy):

My new car uses 2/3 as much gasoline as my old car which used 18 gallons to travel 120 miles how much gasoline does my new car use to travel the same distance?

OpenStudy (diamondboy):

@ Kohai

OpenStudy (diamondboy):

@kohai

OpenStudy (diamondboy):

@SolomonZelman

OpenStudy (solomonzelman):

for old car, \(\large\color{black}{ \displaystyle 18(gallons)=120(miles) }\) and the new car uses only 2/3 of 18 gallons for the same 120 mile distance. for new car, \(\large\color{black}{ \displaystyle (18\times \frac{2}{3})(gallons)=120(miles) }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 18\times \frac{2}{3}=? }\)

OpenStudy (diamondboy):

12

OpenStudy (diamondboy):

gallons

OpenStudy (diamondboy):

thank you very much

OpenStudy (solomonzelman):

yw